OpenStudy (anonymous):

write the points(2,21.75) and (4, 21) as slope intercept form

myininaya (myininaya):

\[m=\frac{21.75-21}{2-4}\]

myininaya (myininaya):

y=mx+b so you got m above now plug in a point (x,y)=(4,21) to find b

OpenStudy (anonymous):

insert a point into the equation y = mx + b and solve for b (x,y)

OpenStudy (anonymous):

slope = (21.75 - 21)/(2-4) = 0.75/-2 = -0.375 now use point-slope y - 21 = -0.375(x - 4) multiply by 1000 1000y - 21000 = -375x + 1500 add 21000 1000y = -375x + 22500 divide by 1000 y = -.375x + 22.5

OpenStudy (anonymous):

with fractions y = (-3/8)x + 45/2